# Driving by Hybrid Navigation: An Online HD-SD Map Association Framework and Benchmark for Autonomous Vehicles

### v2

@@ 6.2 Result
[…]
6.3 Ablation Study
The ablation study experiment is conducted within the OMA-GT dataset, with latency measured using a NVIDIA A6000.
Structure Input Post. Loss Road Pooling Metric
No. PA SA ROPE Boundary Post. CE CTC Avg. Max A-P50:95 La./ms
1 74.1 61
2 62.1 77
3 77.8 64
4 77.9 68
5 78.5 69
6 78.4 70
7 67.7 70
8 78.5 70
9 78.7 70
Table 3: Ablation study of structure, input, loss function and road pooling method. Post. means post process. Avg. means average pooling. La. means latency.
Patch Size 64 256 1024 2048 $\infty$
A-P50:95 77.8 78.4 78.7 78.5 78.4
Latency/ms 77 68 70 72 72
Table 4: Ablation study of patch size of spatial attention.
Group Method N.G Category Path
A-P50:95 78.0 78.2 78.7
Latency/ms 75 73 70
Table 5: Ablation study of group method of path attention. N.G. means non-grouping.
[…]
Visualization and Failed cases. The visualizations and failed cases have been included in the supplemental material. Our approach, as demonstrated by the visualization results, facilitates a stronger long-range correlation on OMA-GT and ensures dependable SD-HD alignment on OMA.
@@ 7 Conclusion
This work introduces OMA, the first dataset for online map association with annotated correspondences with a path-based association Precision-Recall metric aligned with navigation requirements. Furthermore, we introduce MAT, a transformer framework with dual attention mechanisms for map association. MAT achieves 32.3% A-F150:95 / 78.7% A-P50:95 on OMA / OMA-GT with 74 / 70 ms latency, outperforming HMM by 5.5% / 8.6%. The limitation of OMA is its exclusion of dynamic components, such as traffic lights, that affect real-world navigation. In future studies, we plan to incorporate these environmental data into the dataset and assess how it impacts the accuracy of the association.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62433003) and the National Natural Science Foundation of China (62476017).
Appendix A Dataset Analysis
@@
Revised Analysis. As detailed in Section 4.1, the dataset is partitioned into OMA-GT (ground-truth HD maps) and OMA (predicted HD maps), with statistics summarized in Tab. 6. OMA-GT comprises 26,111 training scenarios and 5,613 validation scenarios, totaling 31,724 samples, while OMA contains only 5,573 test scenarios due to the exclusion of low-quality predictions. Both data sets share identical spatial coverage, with HD maps covering (±$(\pm 15m,\pm 30m)$ and SD maps extending to (±$(\pm 75m,\pm 75m)$. Notably, the SD map’s road density in OMA-GT decreases from 15.1 roads/scene during training to 10.8 in validation/test splits, suggesting potential domain shifts between training and evaluation environments.
@@
Quantitative discrepancies between OMA and OMA-GT reveal systemic geometric and topological inconsistencies in predicted maps. OMA predicts an average of 310 lanes/scene, more than four times that of OMA-GT (78.6), with significantly shorter mean lane lengths (2.32 m vs 3.16 m in OMA-GT), indicating both oversegmentation and false positives. This fragmentation is further amplified by OMA’s prediction of 322.06 lane paths/scene (vs. 7.69 in OMA-GT), where ground-truth lanes are frequently split into disconnected fragments. Boundaries exhibit similar degradation: OMA detects 9.65 boundaries/scene (vs 3.40 in OMA-GT) with reduced mean lengths (32.17m vs 44.23m), reflecting fragmented boundary detection. Meanwhile, OMA-GT validation data show slight degradation compared to training splits (e.g. 75.8 vs. 81.3 lanes/scene), highlighting inherent variability in real-world map quality.
@@
Connectivity metrics expose deeper structural errors in OMA predictions. The average lane connectivity in OMA reaches 2.9, substantially higher than OMA-GT’s 2.0/2.1, revealing widespread mislinking of spatially disjoint lanes. Similarly, the OMA-GT validation data show reduced road connectivity (1.8 vs. 2.0 in training), suggesting a domain bias toward simpler topologies in training scenarios. Semantic associations between roads and lanes also degrade significantly. OMA-GT training roads are associated with 1,547 lanes on average, collapsing to 945 in validation splits, which implies degradation of the hierarchical structure in complex scenarios.
@@
These discrepancies have critical implications for benchmarking perception systems. The severe overprediction and fragmentation in OMA highlight the need for metrics penalizing false positives and disconnected paths (e.g., path-length-weighted scores). Furthermore, the mismatch between OMA-GT’s training and validation/test distributions (e.g., road count/length differences) necessitates domain adaptation strategies to ensure generalization. Finally, the collapse of semantic hierarchies in validation data suggests that end-to-end models may struggle to learn robust associations between roads and their constituent elements without explicit structural constraints. Together, these findings underscore the importance of a connectivity-aware association method to avoid overestimating performance on fragmented or mislinked predictions.
@@
Pon split. Both OMA-GT and OMA apply the pon split [27] of the nuScenes dataset [4], ensuring that there is no leakage between the training and validation datasets. For consistent lane prediction segmentation with OMA, we re-trained MapTRv2 [17] using the nuScenes dataset with the pon split. However, this resulted in a significant drop in the quality of the centerline network prediction by MapTRv2 when using the pon split, which adversely affects OMA’s current metrics. Drawing inspiration from the private protocol in MOT17/MOT20 [7], we suggest that future research evaluates the OMA dataset with an enhanced centerline prediction network, without relying on MapTRv2 as a baseline.
Appendix B Visualization
B.1 Attention Map
Figure 7: Visualization of result in OMA-GT.
Figure 8: Visualization of result in OMA.
B.2 Result Compare
Fig.7 illustrates comparisons of model ground truth on OMA-GT, KNN, HMM, MAT and MAT w/ postprocess. The top rows of (d) and (e) exhibit our post-processing module’s effectiveness. MAT predictions initially display incorrect topological connections where roads are mistakenly linked (marked with red circles). Our postprocessing, which utilizes topology-aware beem searching, rectifies this by eliminating non-sequential transitions and reconstructing precise topological paths. The second row demonstrates MAT’s superior handling of complex topologies. Although HMM targets sequential path associations, its single path paradigm often underperforms in complex topologies with intersections. In contrast, our model uses spatial attention to grasp global information and cross-path associations, facilitating adaptive learning of complex topological patterns for accurate connectivity inference. The third row showcases our model’s improved ability to localize associations. Using path-aware attention, the model emphasizes detailed extraction of local features along paths. This targeted local perception ensures precise associations at challenging points, such as junctions, where HMM is typically short due to limited contextual understanding.
Fig. 8 illustrates a comparison of ground truth results for OMA-GT, KNN, HMM, MAT, and MAT with post-processing. Significantly, the visualization demonstrates that our model excels in map association in noisy scenarios with inaccurate centerline predictions, surpassing KNN and HMM by integrating the complementary benefits of global association (SA) and local detail refinement (PA).
Figure 9: Visualization of failed cases.
B.3 Failed Cases
Appendix C Model Families
Methods A-P50 A-P75 A-P95 A-P50:95 La./ms
KNN [5] 72.2 69.9 59.4 68.6 299
HMM [25] 73.8 71.5 60.3 70.1 465
MAT-T 81.2 80.0 68.1 78.2 34
MAT-S 81.4 80.2 68.3 78.3 56
MAT-M 81.5 80.3 68.9 78.6 63
MAT-L (Main Manuscript) 81.6 80.4 69.0 78.7 70
Table 7: Result on OMA-GT. La. means latency.
We develop a family of models derived from those detailed in the main manuscript. By adjusting the number of layers and channels, these models vary in size, memory usage, and inference speed, allowing them to suit various end-side environments. Detailed parameters for each model can be found in Section F of the Supplementary Material.
The experimental results in Tab. 7 demonstrate the effectiveness of our MAT architecture in the accuracy and efficiency dimensions. As the model scales from MAT-T to MAT-L, we observe a consistent improvement in precision metrics (A-P50: 81.2$\rightarrow$81.6, A-P75: 80.0$\rightarrow$80.4, A-P95: 68.1$\rightarrow$69.0) with increasing parameter counts, although at the cost of reduced inference speed (34$\rightarrow$70ms). In particular, even the smallest variant MAT-T outperforms HMM by achieving a 9.6% higher A-P50:95 accuracy while reducing latency by 92.6%, establishing a new Pareto frontier for trajectory association. The minimal performance degradation from A-P50 to A-P95 (-12.6% for MAT-L vs -13.5% for HMM) further highlights the robustness of our dual attention mechanism in preserving topological constraints under stringent matching criteria. The family of MAT allows users to select optimal configurations based on hardware constraints, with potential future extensions toward dynamic computation allocation that activates larger models only during complex topological transitions. These findings quantitatively confirm that explicit modeling of both global topology (SA) and local geometry (PA) through differentiable attention mechanisms can surpass traditional probabilistic approaches while satisfying practical deployment requirements.
Appendix D Metric Details

### v3

@@ 6.3 Ablation Study
Data Efficiency. To investigate the label efficiency of MAT and validate its potential for semi-supervised learning, we conducted experiments using varying subsets of the training data (ranging from 1% to 100%) on MAT-L. The results, summarized in Tab. 6, demonstrate exceptional data efficiency. In particular, with only 5% of the annotated training data, the model achieves a $NR-F1^{50:95}$ of 77.1% in the validation set and 44.5% in the test set. These results are comparable to the performance achieved using 100% of the data (78.7% / 45.0%). This saturation in a low-data regime indicates that MAT effectively captures robust topological features without requiring massive amounts of dense annotations.
Table 6: Data Efficiency experiment on OMA dataset.
Ratio 1% 2% 5% 10% 20% 50% 100%
Val / NR-F150:95 24.9 64.4 77.1 77.7 77.3 78.3 78.7
Test / NR-F150:95 22.3 41.5 44.5 44.6 46.6 44.7 45.0
More Ablation Study, Visualization and Failed cases. Appendix G has shown more study of ablation study of hyperparameters of PA and SA, input size of SD map, beam width of post-process, visualizations, and failed cases.
@@ 7 Conclusion
We propose Online Navigation Refinement (ONR), a new mission that fuses static SD maps with real-time perception for accurate, low-cost lane-level navigation. To achieve ONR, we make three core contributions: We release OMA, the first public benchmark for map associations, and MAT, a transformer-based model for aligning noisy maps. We also introduce Navigation Refinement P-R, a metric evaluating both geometry and association accuracy. For limitations, we will explore end-to-end integration with motion planners, cross-domain adaptation, and unsupervised association learning to reduce annotation dependence.
Ethics Statement
This work uses only publicly available autonomous driving datasets (e.g., nuScenes) and OpenStreetMap (OSM), following their licenses and usage terms. We rely solely on map-level and sensor-derived geometric information and do not use or release any personally identifiable data such as faces, license plates, or raw trajectories linked to individuals. All additional annotations were produced by trained annotators on de-identified map representations. Our method and dataset are intended for research purposes to improve road safety and navigation; any real-world deployment should include thorough testing, safety checks, and compliance with relevant regulations to avoid potential harms from incorrect or misleading guidance.
Reproducibility statement
This work is based entirely on publicly available datasets, specifically the nuScenes autonomous driving dataset and OpenStreetMap (OSM). We describe our model architectures, training objectives, and optimization hyperparameters, as well as dataset preprocessing, train/validation/test splits, and evaluation metrics in the main text and appendix to enable independent verification. Upon publication, we will release our code, including scripts for data preprocessing, map association construction, model training, and evaluation, together with configuration files specifying all hyperparameters, random seeds, and implementation details (e.g., framework and library versions). We will also release the processed map-association annotations introduced in this paper, together with the code of our models and trained checkpoints.
Contents of Appendix
Appendix A LLM Usage
The paper has employed LLM tools such as GPT and Writefull for enhancing and refining writing in the abstract, main text, and appendix.
Appendix B Question and Answer
We have included some Q&A to enable readers to grasp the information in our dataset. Every response follows the double-blind principle.
B.1 Motivation
For what purpose was the dataset created? We introduce this dataset to refine online navigation by transforming road-level paths on SD maps into lane-level paths on online perception maps. This approach facilitates cost-effective and highly real-time lane-level navigation. Current lane-level navigation systems are highly dependent on expensive and globally updated HD maps, which are costly and lag in real-time efficiency. Online perception maps, a significant focus in autonomous driving research, create local HD maps near vehicles using real-time onboard sensor data. However, they lack global topology connections, making them unsuitable for navigation. This dataset envisions aligning SD maps with online perception maps through a map-to-map association paradigm, translating road-level paths on SD maps into lane-level paths on online perception maps.
What is the relationship between this paper and planning methods? The task and methodology of this paper focus on delivering road-level navigation on vehicles and drivers’ devices. Addresses the issue of road-level navigation provision in a context where HD maps are no longer used. Recently, research like NavigScene peng2025navigscene has incorporated SD navigation data into autonomous driving planning modules. We suggest that our approach can act as a preliminary stage for NavigScene, enhancing the SD navigation data to offer lane-level guidance. This grants NavigScene’s planning module more detailed navigation details, thereby increasing planning precision.
B.2 Distribution
Will the dataset be distributed to third parties outside the entity (e.g., company, institution, organization) on behalf of which the dataset was created? Yes, the dataset can be accessed publicly on the Internet.
How will the dataset be distributed (for example, tarball on website, API, GitHub)? The dataset will be released on GitHub and Huggingface.
B.3 Maintenance
Who will be supporting/hosting/maintaining the dataset? The authors will be supporting, hosting, and maintaining the dataset.
How can the owner / curator / manager of the dataset be contacted (e.g., email address)? You can contact the author by email which will be included in the accepted version of the paper.
Is there an erratum? No. We will make a statement if there is any.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Yes. we intend to refresh OMA’s on-line HD map with more robust baselines in the future. The new method will be selected based on comprehensive criteria that include performance measures such as accuracy, efficiency, and generalization.
Will older versions of the dataset continue to be supported/hosted/maintained? Yes.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? Yes. We will make a guide document on the website of OMA.
B.4 Composition
What do the instances that comprise the dataset represent?In OMA, the basic element is a vector that encompasses the road vectors, the lane vectors, and the boundary vectors. Each vector is defined by two points that indicate its position on the map. We have normalized each vector in OMA to match the ego perspective used in real-world perception settings. Furthermore, each vector is equipped with a set of connectivity relationships that specify the other vectors to which it is connected within the road network. Road vectors include an additional road ID to identify their associated road. Lane vectors have been manually marked to show their association with a specific road.
How many instances are there in total (of each type, if appropriate)? OMA contain over 30K scenarios, 480K road path and 2.6M lane vector with manually annotated associations.
Are relationships between individual instances made explicit? On the SD map, all roads, and on the GT OP map, all lanes within the training and validation sets have thorough descriptions and associated annotations. For the test set, we offer descriptions of roads and lanes, but without associated annotations. Nevertheless, the test set can still be evaluated using annotations from the ground truth OP map in the validation set, thanks to the NR P-R metrics introduced in our paper.
Are there recommended data splits (for example, training, development / validation, testing)? Yes. We have already partitioned our dataset into three distinct splits: training, validation, and testing.
B.5 Collection Process
Who was involved in the data collection process (e.g., students, crowd workers, contractors)? The annotations are provided by experienced annotators and multiple validation stages.
B.6 Use
What (other) tasks could the dataset be used for? This dataset is primarily used for map association, a new task for online navigation refinement.
Figure 6: Challenge for map association. (a) GPS shift, (b) Semantic disparities, (c) missing lane, (d) noise in OP map
Appendix C Related Work
C.1 End-to-end Autonomous Driving
Planning, being an essential component for numerous fields zhang2025investigating; zhang2024exploring; zhang2025fully; bai2024learning, is recognized as one of the pivotal research domains in the realm of autonomous driving at present. VAD jiang2023vad utilizes an ego query mechanism to forecast individual mode trajectories, whereas VADv2 chen2024vadv2 enhances this by adopting a probabilistic framework that considers multiple trajectories. SparseDrive sun2025sparsedrive is innovative in its creation of parallel motion and planning modules, effectively decreasing the computational requirements of the BEV features. DiffusionDrive liao2025diffusiondrive employs a truncated diffusion policy to improve the probabilistic depiction of trajectories, while MomAD song2025don aims to enhance stability and maintain consistency throughout sequential planning decisions. Significantly, unlike other end-to-end planning techniques, NavigScene peng2025navigscene uses navigation data based on SD maps, achieving more accurate and reliable autonomous driving, thus underscoring the significance of navigation data. Consequently, we propose that the refinement of lane-level navigation through MAT could further enhance the capabilities of current autonomous driving methodologies.
C.2 Mission Comparison
Fig. 7 compares four tasks: SD map navigation, map generation, map matching, and map association. Initially, navigation based on the standard definition (SD) map achieved a road-level path, which is now widely integrated into navigation apps. However, road-level navigation often does not provide precise directions for specific lanes and directions. As a result, the development of economical and quick lane-level navigation systems has emerged as a significant area of GIS. Next, lane-level online perception (OP) maps are generated using data from vehicle sensors by map generation. Unfortunately, such maps do not have a global topological framework, making it challenging to form a route from the start of the navigation to the endpoint. Subsequently, the task of map matching attempts to align GPS path with global SD or HD maps by following a path-to-map matching model. However, this model falls short when trying to match SD and OP maps due to disregarding the differences between these types. Finally, the map association task seeks to create linkages between SD and OP maps via a map-to-map association model. Using the SD map, the map association identifies a global route from the origin to the destination. Then this global route is meticulously translated into the OP map at the lane level through the association of the SD and OP maps, resulting in a comprehensive lane-level navigation path.
Figure 7: The comparison of (a) SD map navigation, (b) mapping generation, (c) map matching and (d) map association. The blue word is the input of the task and the purple word is the output of the task
Appendix D Dataset
Table 7: Statistics of OMA train, val and test set.
   Split    Train    Val    Test
   HD map Range    $(\pm 15m,\pm 30m)$
   SD map Range    $(\pm 75m,\pm 75m)$
   Scene Segment    26111    5613    5573
   Avg. lane per scene    81.3    75.8    310.34
   Avg. lane path per scene    7.40    7.97    322.06
   Avg. boundary per scene    3.48    3.31    9.65
   Avg. length per lane    3.14m    3.19m    2.32m
   Avg. length per boundary    44.81m    43.64m    32.17m
   Avg. road per scene    15.1    10.8    10.8
   Avg. length per road    38.2m    50.3m    50.3m
   Avg. Connection per lane    2.0    2.1    2.9
   Avg. Connection per road    2.0    1.8    1.8
   Avg. Connection per boundary    2.0    2.0    2.0
D.1 Revised Analysis.
@@ Appendix D Dataset
As detailed in Section 4.1, the dataset is partitioned into OMA train, val, and test set, with statistics summarized in Tab. 7. The OMA training and val set comprises 26,111 training scenarios and 5,613 validation scenarios, totaling 31,724 samples, while the test set contains only 5,573 test scenarios due to the exclusion of low-quality predictions. Both data sets share identical spatial coverage, with HD maps covering $(\pm 15m,\pm 30m)$ and SD maps extending to $(\pm 75m,\pm 75m)$. Notably, the SD map’s road density in OMA train and val set decreases from 15.1 roads/scene during training to 10.8 in validation/test splits, suggesting potential domain shifts between training and evaluation environments.
@@
Quantitative discrepancies between train, val and test set reveal systemic geometric and topological inconsistencies in predicted maps. The test set predicts an average of 310 lanes/scene, more than four times that of the Train and Val set (73.8), with significantly shorter mean lane lengths (2.32 m vs 3.16 m in trainval), indicating both over-segmentation and false positives. This fragmentation is further amplified by test’s prediction of 322.06 lane paths/scene (vs. 7.69 in Train), where ground-truth lanes are frequently split into disconnected fragments. Boundaries exhibit similar degradation: test set detects 9.65 boundaries/scene (vs 3.40 in trainval) with reduced mean lengths (32.17m vs 44.23m), reflecting fragmented boundary detection. Meanwhile, validation data show slight degradation compared to training splits (e.g. 75.8 vs. 81.3 lanes/scene), highlighting inherent variability in real-world map quality.
@@
Connectivity metrics expose deeper structural errors in test set predictions. The average lane connectivity in test set reaches 2.9, substantially higher than trainval’s 2.0/2.1, revealing widespread mislinking of spatially disjoint lanes. Similarly, the validation data show reduced road connectivity (1.8 vs. 2.0 in training), suggesting a domain bias toward simpler topologies in training scenarios. Semantic associations between roads and lanes also degrade significantly. Training roads are associated with 1,547 lanes on average, collapsing to 945 in validation splits, which implies degradation of the hierarchical structure in complex scenarios.
@@
These discrepancies have critical implications for benchmarking perception systems. The severe over-prediction and fragmentation in test set highlight the need for metrics penalizing false positives and disconnected paths (e.g., path-length-weighted scores). Furthermore, the mismatch between training and validation/test distributions (e.g., road count/length differences) necessitates domain adaptation strategies to ensure generalization. Finally, the collapse of semantic hierarchies in validation data suggests that end-to-end models may struggle to learn robust associations between roads and their constituent elements without explicit structural constraints. Together, these findings underscore the importance of a connectivity-aware association method to avoid overestimating performance on fragmented or mislinked predictions.
Table 8: Refinement statistics for different scenes in OMA
Scene name Number of centerlines Number of refinements Refinement ratio
Boston 1,205,661 2,341 0.194%
Singapore 910,140 3,682 0.404%
D.2 Range of Sample
The model is trained on nuScenes using synchronized LiDAR and camera inputs with official configuration. For each sample, an SD map cropping measure $150\,m\times 150\,m$ centered around the ego vehicle preserves the adjacent topological context followed by the sensor setting of nuScenes caesar2020nuscenes. For the OP map, the cropping measure of the OP map $30\,m\times 60\,m$ was centered around the vehicle of the ego, as referenced in maptrv2; liu2023vectormapnet; li2022hdmapnet.
D.3 Pon split and OP maps
@@
Both Train, val and test set apply the pon split roddick2020predicting of the nuScenes dataset nuscenes2019, ensuring that there is no leakage between the training and validation datasets. For consistent lane prediction segmentation with OMA test set, we re-trained MapTRv2 maptrv2 using the nuScenes dataset with the pon split. Drawing inspiration from the private protocol in MOT17/MOT20 dendorfer2021motchallenge, we suggest that future research evaluates the test set with an enhanced centerline prediction network, without relying on MapTRv2 as a baseline. Furthermore, we will establish an open evaluation protocol allowing submissions to include their own OP maps, thereby reducing reliance on a single model.
D.4 Geographical generalization
Regarding the geographical generalization of OMA, we elaborate from two aspects: First, OMA’s data originates from the nuScenes dataset, which is widely used in the autonomous driving field and contains data from two different countries, Singapore and Boston. Given that nuScenes has become a core dataset in the autonomous driving field for perception, mapping, and planning since its introduction, we believe that the nuScenes data itself possesses certain representativeness and generalizability. Second, we conducted cross-validation ablation experiments in the main text. The cross-experimental results of MAT on OMA demonstrate that MAT models trained on different geographical regions possess certain cross-regional generalization capabilities, reflecting the good geographical generalizability of the OMA dataset itself.
D.5 Annotation Workflow
The OMA data annotation process is divided into two primary stages: First, we performed coarse alignment based on the GPS coordinates from the SD map of OSM and the GT OP map of NuScenes. Second, we employed a skilled annotation team to correlate the SD map with the real perception map. The specialists were then engaged to evaluate and improve the annotations, ensuring their ultimate quality. Tab. 8 presents the total count and percentage of data changes in various regions in Phase 3. It is evident that the alteration rate for Singapore and Boston remained less than 1%, suggesting a generally high standard of data annotation quality.
D.6 Handling Transition Ambiguity
Addressing the concern regarding transition lines, we acknowledge that assigning roads in such transition areas presents inherent ambiguity due to the heterogeneity between SD and OP Maps. To mitigate potential impacts on training stability and evaluation fairness, we implemented specific strategies: (1) Standardized Annotation Protocol: We established a unified rule where the assignment of a Lane Vector at an SD Link transition boundary is determined by the SD Link closest to the Lane Vector’s start point, ensuring topological consistency. (2) Tolerance in Evaluation Metrics: Our metric assesses the precision of path-level associations across 11 thresholds (50% to 95%). Crucially, the 5% tolerance buffer included even at the strictest threshold is specifically designed to accommodate inevitable semantic ambiguity at transition boundaries.
D.7 Future Expansion Strategy
We acknowledge the importance of scaling the benchmark to encompass more diverse datasets (e.g., Argoverse). To minimize annotation costs during future expansion, we propose a “Human-in-the-Loop” iterative annotation strategy. This approach is strongly supported by our data efficiency experiments as shown in Tab. 6, which demonstrated that MAT can achieve competitive performance with as little as 5% of the training data. The specific pipeline is as follows:
  1. 1.
    Model-Assisted Pre-annotation: We utilize the MAT model trained on the existing OMA (nuScenes) dataset to perform zero-shot or few-shot inference on unlabeled new data. This generates initial “draft” associations (e.g., SD-to-HD correspondences and topology) and significantly reduces the cold-start problem.
  2. 2.
    Lightweight Manual Correction: Instead of labeling from scratch, annotators focus solely on verifying and correcting the model’s high-confidence predictions. Given the model’s high data efficiency, the pre-annotation quality improves rapidly even with a small set of initial corrections.
  3. 3.
    Closed-Loop Iteration: The corrected data is immediately integrated into the training set to fine-tune the model. The updated model is then used to pre-annotate the subsequent batch of data, creating a positive feedback loop that progressively reduces manual workload.
Appendix E Metric
E.1 Details of NR-PR